MID-INFRARED SPECTRUM OF THE CO_{2}-SO_{2} VAN DER WAALS COMPLEX

\underline{HANS D. OSTHOFF} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2, Canada.

We have recorded the mid-infrared spectrum of the CO_{2}-SO_{2} van der Waals complex near the asymmetric stretching frequency of the R(0) transition of the CO_{2} monomer at 2349.917 cm^{−1}. The spectra were recorded on our new molecular beam tunable diode laser spectrometer, in which the sample is injected axially to the main optical axis of the Herriott multipass cell, resulting in an increased residence time of the molecules in the path of the laser beam. The assigned transitions in the mid-infrared spectrum of CO_{2}-SO_{2} are c-type transitions in agreement with spin statistical analysis. Transitions originating from levels with odd K_{a} in the vibrational ground state are absent from the spectrum. The assignments of several transitions was directly confirmed by microwave-infrared (MW-IR) double resonance experiments, using several known^{a} a-type microwave transitions of the complex. The main spectroscopic constants are ν_{0}=2349.6160(11) cm^{−1}, A”=5544.81(40) MHz, B”=1510.034(37) MHz, C”=1489.402(36) MHz, D_{J}”=4.578(16) kHz, A’=5524.34(63) MHz, B’=1507.988(164) MHz, C’=1488.404(214) MHz, and D_{J}’=4.63(80) kHz. The assignment of additional observed transitions is in progress.

aL. Sun, I. I. Ioannou and R. L. Kuczkowski Mol. Phys. 88, 255-268 (1996).